# An Experimental Design Framework for Label-Efficient Supervised Finetuning of Large Language Models

### v1

@@ Abstract
Supervised finetuning (SFT) on instruction datasets has played a crucial role in achieving the remarkable zero-shot generalization capabilities observed in modern large language models (LLMs). However, the annotation efforts required to produce high quality responses for instructions are becoming prohibitively expensive, especially as the number of tasks spanned by instruction datasets continues to increase. Active learning is effective in identifying useful subsets of samples to annotate from an unlabeled pool, but its high computational cost remains a barrier to its widespread applicability in the context of LLMs. To mitigate the annotation cost of SFT and circumvent the computational bottlenecks of active learning, we propose using experimental design. Experimental design techniques select the most informative samples to label, and typically maximize some notion of uncertainty and/or diversity. In our work, we implement a framework that evaluates several existing and novel experimental design techniques and find that these methods consistently yield significant gains in label efficiency with little computational overhead. On generative tasks, our methods achieve the same generalization performance with only $50\%$ of annotation cost required by random sampling.
Figure 1: Comparison between different annotation schemes for label-efficient SFT. Random sampling simply chooses prompts uniformly at random, but one can choose them more strategically both through active learning and experimental design. Active learning, however, is an adaptive procedure and requires computationally expensive model retraining and inference for every batch of annotation. In this paper, we study the problem through the lens of experimental design, which enjoys increasing label-efficiency compared to random sampling, while incurring minimal computation cost compared to active learning.
@@ 1 Introduction
Supervised finetuning (SFT) on instruction datasets has shown immense potential in improving the zero-shot performance of large language models (LLMs) (Wei et al., 2022). Recent developments in this field have been fostered by the availability of large-scale instruction datasets, consisting of natural language instructions with desired responses based on human judgment (Wei et al., 2022; Mishra et al., 2022; Longpre et al., 2023). Throughout the community, there have been several efforts to further increase the number of tasks included in these datasets to improve LLM generalization (Xu et al., 2022; Wang et al., 2022; Honovich et al., 2022; Wang et al., 2023c).
Scaling the size of these datasets requires annotating a massive set of instructions with detailed responses, which can be done either manually or automatically. Manual approaches, through the use of crowd-workers or expert annotators, can produce high quality responses for almost any task. However, producing human labels for a full dataset can become extremely expensive at scale due to the labor intensive process of annotation. More recently, automatic annotation methods have been proposed to reduce the burden on human annotators by labeling instructions using preexisting LLMs (Honovich et al., 2022; Wang et al., 2023c; Peng et al., 2023). However, such techniques are not universally applicable since existing LLMs may not be able to produce high quality annotations for domain-specific tasks and datasets. Furthermore, models like GPT-4 are costly to query, so generating responses for a large set of prompts may be prohibitive. Therefore, we seek to answer the following question: Can LLMs generalize well when finetuned on fewer annotated prompts?
@@
Overall, our contributions can be summarized as follows: (1) we are the first, to the best of our knowledge, to utilize experimental design for SFT (2) we introduce a framework to perform comprehensive evaluation on existing experimental design techniques (3) we propose a suite of novel strategies that improve the label efficiency, significantly outperforming random sampling by more than $2\%$ accuracy across different annotation budgets (4) compared to previous works (Kung et al., 2023; Perlitz et al., 2023), our work is the first to see annotation cost savings on generative tasks by using only $50\%$ of annotation budget (compared to random sampling) to achieve the same generalization performance.
@@ 2 Related Work
Our study focuses on experimental design, rather than active learning for two main reasons. First, active learning generally requires retraining the model after each batch of labels is gathered in order to select the next batch for labeling, or requires some other computationally expensive process to leverage the gathered information. Experimental design does not require additional computation of this sort, since the subset selection of unlabeled examples is done just once before any labels are collected. This also allows for parallelization, in the sense the the labeling tasks can be conducted separately from each other. The second reason for our focus on experimental design is that under fairly mild assumptions, it is mathematical known that active learning can not provide a significant gain (more than a factor of two) over optimal experimental design. This result comes from the theory of information-based complexity (IBC); see Theorem 5.2.1 in Traub et al. (1988). At a high level, the IBC theory shows that if the model class is convex and symmetric and the task is to minimize a regression loss (e.g., squared error as opposed to a 0/1 classification loss), then the expected loss resulting from an optimal experimental design is no more than twice that of the expected loss using active learning, in general. The convexity of the model class is crucial for this result. Strict classifiers generally do not form a convex class (e.g., convex combinations of separating hyperplanes are not separating hyperplanes, but convex combinations of logistic regression functions are regression functions). As LLMs care about the entire softmax distribution over high likelihood tokens, we argue that active learning is unlikely to provide significant gains over good experimental designs in fine-tuning LLMs.
@@
Data selection strategy for label-efficient learning has been largely studied under the framework of deep active learning, which sequentially and adaptively chooses informative examples to annotate. Deep active learning methods typically use measures of uncertainty (Atlas et al., 1989; Settles, 2011; Gal et al., 2017; Ducoffe and Precioso, 2018; Beluch et al., 2018), diversity (Sener and Savarese, 2018; Geifman and El-Yaniv, 2017; Citovsky et al., 2021), or some combination of both (Wei et al., 2015; Ash et al., 2019, 2021; Zhang et al., 2022; Nuggehalli et al., 2023) in order to determine a set of samples to annotate. In addition to classical instance-level selection, there also many existing works focused on task-wise selection Xu et al. (2023); Chen et al. (2023b); Wang et al. (2023a); Fifty et al. (2021). While we propose an experimental design framework for the label-efficient SFT problem, our strategies are inspired by the latest instance-level deep active learning literature and adapted to work under this framework for supervised finetuing of LLMs.
A few works have considered the application of active learning for tasks related to SFT. Perlitz et al. (2023) explores the use of active learning to improve the label efficiency of the closely related task of natural language generation. However, they report inconsistent findings showing that there are limited settings where active learning demonstrates advantages over random sampling. Kung et al. (2023) is the most similar to our work, proposing an active learning framework for instruction tuning. Unlike our work, Kung et al. (2023) performs task-level selection, as opposed to instance-level selection. In other words, their framework estimates the usefulness of each unlabeled task and annotates all instances within the tasks that are deemed most useful. However, task-level selection makes the simplifying assumption that every instance is equally useful within a task, which may inhibit the quality of the selected subset. Moreover, the approach in Kung et al. (2023) suffers from the aforementioned computational cost of active learning which is avoided in our framework with the use of experimental design.
@@
Many recent works demonstrate that small subsets of instruction data can be sufficient for finetuning an LLM. Zhou et al. (2023) demonstrate that finetuning an LLM on a dataset consisting of only manually-curated 1000 instruction/response pairs is sufficient to achieve strong generalization, but do not propose any general algorithmic procedure for subset selection. Other works propose instruction dataset pruning techniques that select subsets based on some combination of quality, diversity, and/or difficulty (Chen et al., 2023a; Bukharin and Zhao, 2023; Du et al., 2023; Li et al., 2023). These approaches use both the instruction and its corresponding response to choose which training samples should be retained to improve the computational efficiency of SFT. Unlike these approaches, our framework is designed to maximize label efficiency and assumes that the response to an instruction are not available until selected for annotation.
@@ 3 Prompt Selection Strategy
Under the experimental design framework, the learner is given a set of initial $N$ prompts $X=\{x_{1},x_{2},...,x_{N}\}$, where each prompt $x_{i}$ is a sequence of input tokens of length $l$, $x_{i}=\{x_{i,1},...,x_{i,l}\}$. Additionally, we let $g$ denote the pretrained language model. Given a budget of $k<N$, a selection strategy chooses $k$ prompts from $X$, denoted as $S\subset X$, based on different measures of informativeness of annotating an example to the pretrained model $g$. One then gathers well-written responses to prompts in $S$ from annotators (e.g., human experts or advanced LLMs). Experimental design aims to optimize the performance of the model $g^{\prime}$ finetuned on the selected prompts $S$ and their responses.
@@ 3.1 Uncertainty-Based Selection
Below, we provide four instances of uncertainty measures. Formally, we let the pretrained model (with greedy decoding) $g:X\rightarrow(\mathcal{D}\times\Delta_{\mathcal{D}})^{[L]}$ map prompts to a sequence of tokens and softmax scores up to length $L$. Here each token is from a dictionary $\mathcal{D}$ and each corresponding softmax scores lives in the probability simplex $\Delta_{\mathcal{D}}$ over the dictionary. We let $g_{y}(\cdot)$ denote the sequence of tokens and $g_{p}(\cdot)$ denote the sequence of softmax scores.
Mean Entropy (Settles, 2011; Kremer et al., 2014) measures the tokenwise negative entropy of the softmax probability scores. The uncertainty measure is taken as the mean across tokens as follows:
@@
Least Confidence (Settles, 2009, 2011) measures the model’s confidence as the product of probabilities of the generated sequence . A model is more confident when the likelihood of the generated sequence is high. We take the negative confidence score as the uncertainty, which allows us to choose the least confident sequences for annotation:
@@
Mean Margin (Tong and Koller, 2001; Balcan et al., 2006; Settles, 2011) measures uncertainty by taking the different between most likely and second likely token for each element in the generated sequence . A higher difference corresponds with a clearer separation between the model’s best choice from its second best choice. We use the negative margin score average over all tokens as the uncertainty measure:
@@
Min Margin is a novel strategy where we measure uncertainty based on the token with the smallest margin score instead of taking the average over all tokens. Intuitively, two sequences could have equal average token-wise margin score, but the one with smaller minimum margin is more likely to be generated as a different sequence with top-2 decoding. Mathematically, this is defined as:
$\widetilde{U}_{\text{margin}}(x)=-\left(\min_{p\in g_{p}(x)}\beta_{1}(p)-\beta_{2}(p)\right).$
@@
where $f$ is a feature extractor mapping prompts into feature space in $\mathbb{R}^{d}$ and is derived from the pretrained model $h$. For decoder-only architectures, we use the first hidden state as the feature. To optimize the above NP-hard object (Cook et al., 1998), we follow the greedy methods proposed by Sener and Savarese (2018), which enjoys a 2 multiplicative approximation guarantee to the optimal selection.
@@ 3.3 Submodular Selection
We primarily use the radial basis function as a similarity metric, where $w_{ij}=\exp{(-\frac{\|f(x_{i})-f(x_{j})\|^{2}}{\gamma})}$ and $\gamma>0$ is a controllable hyperparameter often referred to as the kernel width. Intuitively, $\gamma$ tunes is the degree of similarity between two data points. As $\gamma$ decreases, the similarity between $x_{i}$ and $x_{j}$ also decreases. In limiting case as $\gamma\to 0$, similarity $w_{ij}=0$ for all $i\neq j$. For situations where hyperparameter tuning is not feasible, we propose an alternate similarity metric, where $w_{ij}=\max\{0,\frac{f(x_{i})^{T}f(x_{j})}{\|f(x_{i})\|\|f(x_{j})\|}\}$.
@@
Table 1: Massive Multitask Language Understanding (MMLU) evaluation of strategies under different annotation budgets. Each result of random strategy is averaged over 6 seeds due to the high variance from both data selection and training. Other results are averaged over 3 random seeds where the randomness mainly comes from the training. The confidence intervals are based on standard error.
@@
Table 2: Big-Bench-Hard chain-of-thoughts (BBH-CoT) evaluation of strategies under different annotation budgets. Each result of random strategy is averaged over 6 seeds where each result of other strategies is averaged over 3 random seeds with standard error shown as confidence interval.
@@ 4.1 Experiment Setup
FLAN V2 (Longpre et al., 2023) is a widely-used instruction fine-tuning dataset that combines FLAN 2021, P3++, Super-Natural Instructions, along with additional reasoning, dialogue, and program synthesis datasets. We utilize a 100K subset of FLAN V2, processed by (Wang et al., 2023b), as our training data pool, from which we select prompts and annotations (i.e., responses).
@@
We conduct experiments with the 7B version of the prefix language model LAMMA-2 (Touvron et al., 2023), across different annotation budgets. Prior to fine-tuning, we choose a subset of prompts for annotation by either random sampling or using experimental design strategies, including uncertainty-based selection, k-Center selection, and submodular selections. These strategies are computed based on the prefix model only. Subsequently, we fine-tune the model on the annotated prompt/response pairs using Low-rank Adaptation (LoRA) (Hu et al., 2021).
@@
We adopt the similar evaluation tasks as in the original FLAN V2, using the MMLU (Hendrycks et al., 2020) and BBH (Suzgun et al., 2022) benchmarks to evaluate the zero-shot generalization capability of our fine-tuned model. Massive Multitask Language Understanding dataset (MMLU) is a classification task with a set of questions about 57 subjects ranging in difficulty from elementary levels to professional levels, broadly testing mode’s factual knowledge and reasoning. Big-Bench-Hard (BBH) is a generation task with 23 challenging tasks from Big-Bench (Srivastava et al., 2022), broadly testing models’ general reasoning capabilities. In our study, we specifically select 5-shot MMLU and a random 20% subset of Chain-Of-Thought BBH inputs, due to the computational source limitations. While more evaluation metrics can be employed, we believe these two are sufficient to demonstrate preliminary findings.

### v3

@@ Abstract
Supervised finetuning (SFT) on instruction datasets has played a crucial role in achieving the remarkable zero-shot generalization capabilities observed in modern large language models (LLMs). However, the annotation efforts required to produce high quality responses for instructions are becoming prohibitively expensive, especially as the number of tasks spanned by instruction datasets continues to increase. Active learning is effective in identifying useful subsets of samples to annotate from an unlabeled pool, but its high computational cost remains a barrier to its widespread applicability in the context of LLMs. To mitigate the annotation cost of SFT and circumvent the computational bottlenecks of active learning, we propose using experimental design. Experimental design techniques select the most informative samples to label, and typically maximize some notion of uncertainty and/or diversity. In our work, we implement a framework that evaluates several existing and novel experimental design techniques and find that these methods consistently yield significant gains in label efficiency with little computational overhead. On generative tasks, to reach the same generalization performance, our methods save $50\%$ of the annotation cost compared to random sampling.
Figure 1: Comparison between different annotation schemes for label-efficient SFT. Random sampling simply chooses prompts uniformly at random which underperforms as it is prone to redundancy and may oversample from the major modes. On the other hand, one can choose them more strategically both through active learning and experimental design. Active learning, however, is an adaptive procedure and requires computationally expensive model retraining and inference for every batch of annotation. In this paper, we study the problem through the lens of experimental design, which enjoys increased label-efficiency compared to random sampling, while incurring minimal computation cost compared to active learning.
@@ 1 Introduction
Supervised finetuning (SFT) on instruction datasets has shown immense potential in improving the zero-shot performance of large language models (LLMs) (Wei et al., 2022). Recent developments in this field have been fostered by the availability of large-scale instruction datasets, consisting of natural language instructions with desired responses based on human judgment (Wei et al., 2022; Mishra et al., 2022; Longpre et al., 2023). Throughout the community, there have been several efforts to further increase the number of tasks included in these datasets to improve LLM generalization (Xu et al., 2022; Wang et al., 2022; Honovich et al., 2022; Wang et al., 2023c). In addition, supervised finetuning is especially important for handling novel forms of tasks. For example, to develop new multi-modal capabilities or defend against novel prompt hacking strategies, one must need human generated responses to finetune existing state-of-art models like GPT-4.
A crucial bottleneck of SFT is the need for annotating a massive set of instructions with detailed responses, which can be done either manually or automatically. For improving state-of-art models, one necessarily need to rely on manual approaches. This involves the use of crowd-workers or expert annotators, who produce high quality responses for almost any task but can become extremely expensive at scale due to the labor intensive process of annotation. To improve small scale and more domain-specific language models, automatic annotation methods have been proposed to reduce the burden on human annotators by labeling instructions using preexisting LLMs (Honovich et al., 2022; Wang et al., 2023c; Peng et al., 2023). However, models like GPT-4 are still costly to query, making the annotation for a large set of prompts potentially prohibitive. In addition, for specialized domains, human experts are still crucial in this annotation process as general LLMs may not provide high quality responses. Therefore, we seek to answer the following question:
Can we propose label-efficient strategies that annotate fewer prompts while obtaining equally good generalization performance?
@@
Overall, our contributions can be summarized as follows: (1) we are the first, to the best of our knowledge, to utilize experimental design for SFT (2) we introduce a framework to perform comprehensive evaluation on existing experimental design techniques (3) we propose a suite of novel strategies that improve the label efficiency, significantly outperforming random sampling by more than $2\%$ accuracy across different annotation budgets (4) compared to previous works (Kung et al., 2023; Perlitz et al., 2023), our work is the first to see annotation cost savings on generative tasks. To reach the same generalization performance, our methods save $50\%$ of the annotation cost compared to random sampling (unlike random sampling that fails to achieve the same generalization).
@@ 2 Related Work
Our study focuses on experimental design, rather than active learning. Active learning generally requires retraining the model and forward inference on the entire pool after each batch of labels is gathered in order to select the next batch for labeling. The retraining and repeated inference can be computationally expensive, particularly in the case of LLMs. Experimental design does not require additional computation of this sort, since the subset selection of unlabeled examples is done just once before any labels are collected.
@@
Data selection strategies for label-efficient learning have been largely studied under the framework of deep active learning, which sequentially and adaptively choose informative examples to annotate. Deep active learning methods typically use measures of uncertainty (Atlas et al., 1989; Settles, 2011; Gal et al., 2017; Ducoffe and Precioso, 2018; Beluch et al., 2018), diversity (Sener and Savarese, 2018; Geifman and El-Yaniv, 2017; Citovsky et al., 2021), or some combination of both (Wei et al., 2015; Ash et al., 2019, 2021; Zhang et al., 2022; Nuggehalli et al., 2023) in order to determine a set of useful samples to annotate. In addition to classical instance-level selection, there are also many existing works focused on task-wise selection Xu et al. (2023); Chen et al. (2023b); Wang et al. (2023a); Fifty et al. (2021). While we propose an experimental design framework for the label-efficient SFT problem, our strategies are inspired by the latest instance-level deep active learning literature and adapted to work under this framework for supervised finetuning of LLMs.
A few works have considered the application of active learning for SFT or other closely related settings. Perlitz et al. (2023) explores the use of active learning to improve the label efficiency of the closely related task of natural language generation. However, this work reports inconsistent findings showing that there are limited settings where active learning demonstrates any significant advantages over random sampling. Kung et al. (2023) is the most similar to our work, proposing an active learning framework for instruction tuning. Unlike our work, Kung et al. (2023) performs task-level selection, as opposed to instance-level selection. In other words, their framework estimates the usefulness of each unlabeled task and annotates all instances within the tasks that are deemed most useful. However, task-level selection makes the simplifying assumption that every instance is equally useful within a task, which may inhibit the quality and reduce the resolution of the selected subset. Moreover, the approach in Kung et al. (2023) suffers from the aforementioned computational cost of active learning which is avoided in our framework with the use of experimental design.
@@
Many recent works demonstrate that small subsets of instruction data can be sufficient for finetuning an LLM. Zhou et al. (2023) demonstrate that finetuning an LLM on a dataset consisting of only manually-curated 1000 instruction/response pairs is sufficient to achieve strong generalization, but do not propose any general algorithmic procedure for subset selection. Other works propose instruction dataset pruning techniques that select subsets based on some combination of quality, diversity, and/or difficulty (Chen et al., 2023a; Bukharin and Zhao, 2023; Du et al., 2023; Li et al., 2023a, 2024). These approaches use both the instruction and its corresponding response to choose which training samples should be retained to improve the computational efficiency of SFT. Unlike these approaches, our framework is designed to maximize label efficiency and assumes that the response to an instruction are not available until selected for annotation. Finally, Hu et al. (2023) propose techniques to reduce the annotation cost to construct validation sets that are used for model selection, which is complementary to the method proposed in our work.
@@ 3 Prompt Selection Strategy
Under the experimental design framework, the learner is given a set of initial $N$ prompts $X=\{x_{1},x_{2},...,x_{N}\}$, where each prompt $x_{i}\in\mathcal{X}$ is a sequence of input tokens of length $\ell_{i}$, $x_{i}=\{x_{i,1},...,x_{i,\ell_{i}}\}$, where $\mathcal{X}$ denotes the domain of all possible sequence of input tokens. Additionally, we let $g$ denote the pretrained language model. Given a budget of $k<N$, a selection strategy chooses $k$ prompts from $X$, denoted as $S\subset X$, based on different measures of informativeness of annotating an example to the pretrained model $g$. One then gathers well-written responses to prompts in $S$ from annotators (e.g., human experts or advanced LLMs). Experimental design aims to optimize the performance of the model $g^{\prime}$ finetuned on the selected prompts $S$ and their responses.
@@ 3.1 Uncertainty-Based Selection
Below, we provide four instances of uncertainty measures. Formally, we let the pretrained model (with greedy decoding) $g:\mathcal{X}\rightarrow(\mathcal{D}\times\triangle_{\mathcal{D}})^{[L]}$ map prompts to a sequence of predicted tokens and the softmax probability distribution at each decoding step, up to length $L$. Here each token is from a dictionary $\mathcal{D}$ and each corresponding softmax probability distribution lives in the probability simplex $\triangle_{\mathcal{D}}$ over the dictionary. We let $g_{y}(\cdot)$ denote the sequence of $L$ tokens and $g_{p}(\cdot)$ denote the sequence of $L$ softmax probability distributions.
Mean Entropy
(Settles, 2011; Kremer et al., 2014) measures the tokenwise negative entropy of the softmax probability scores. The uncertainty measure is taken as the mean across tokens as follows:
@@
Where $p_{t}$ represents the softmax score (i.e., the probability) of the token $t$, and the entire set of $p$ values represents the distribution across the entire dictionary for a specific position in an $L$-word sentence.
Least Confidence
(Settles, 2009, 2011) measures the model’s confidence as the product of probabilities of the generated sequence. A model is more confident when the likelihood of the generated sequence is high. We take the negative confidence score as the uncertainty, which allows us to choose the least confident sequences for annotation:
@@
Mean Margin
(Tong and Koller, 2001; Balcan et al., 2006; Settles, 2011) measures uncertainty by taking the different between most likely and second likely token for each element in the generated sequence . A higher difference corresponds with a clearer separation between the model’s best choice from its second best choice. We use the negative margin score average over all tokens as the uncertainty measure:
@@
Min Margin
is a novel strategy where we measure uncertainty based on the token with the smallest margin score instead of taking the average over all tokens. Intuitively, two sequences could have equal average token-wise margin score, but the one with smaller minimum margin is more likely to be generated as a different sequence with top-2 decoding. Mathematically, this is defined as:
$\widetilde{U}_{\text{min margin}}(x)=-\left(\min_{p\in g_{p}(x)}\beta_{1}(p)-\beta_{2}(p)\right).$
@@
where $f$ is a feature extractor mapping prompts into feature space in $\mathbb{R}^{d}$ and is derived from the pretrained model $h$. For decoder-only architectures, we use the last hidden state as the feature. To optimize the above NP-hard object (Cook et al., 1998), we follow the greedy methods proposed by Sener and Savarese (2018), which enjoys a 2 multiplicative approximation guarantee to the optimal selection.
@@ 3.3 Submodular Selection
We primarily use the radial basis function as a similarity metric, where $w_{ij}=\exp{(-\frac{\|f(x_{i})-f(x_{j})\|^{2}}{\gamma})}$ and $\gamma>0$ is a controllable hyperparameter often referred to as the kernel width. Intuitively, $\gamma$ tunes is the degree of similarity between two data points. As $\gamma$ decreases, the similarity between $x_{i}$ and $x_{j}$ also decreases. In limiting case as $\gamma\to 0$, similarity $w_{ij}=0$ for all $i\neq j$. For situations where hyperparameter tuning is not feasible, we propose an alternate function where the similarity metric is fixed and not tuned, where $w_{ij}=\max\{0,\frac{f(x_{i})^{T}f(x_{j})}{\|f(x_{i})\|\|f(x_{j})\|}\}$.
@@
Table 1: Massive Multitask Language Understanding (MMLU) evaluation of models trained on subsets selected by strategies from a pool of 99k under different annotation budgets. Each result of random strategy is averaged over 6 seeds due to the high variance from both data selection and training. Other results are averaged over 3 random seeds where the randomness mainly comes from the training. The confidence intervals are based on standard error.
@@
Table 2: Big-Bench-Hard chain-of-thoughts (BBH-CoT) evaluation of models trained on subsets selected by strategies from a pool of 99k under different annotation budgets. Each result of random strategy is averaged over 6 seeds where each result of other strategies is averaged over 3 random seeds with standard error shown as confidence interval.
@@ 4 Experiments
In this section, we compare multiple experimental design strategies against random sampling and observe significant improvements. We describe the various setups in Section 4.1 and report the evaluation results based on common benchmarks (Section 4.2) and comparisons by GPT-4 (Section 4.3). Lastly, we conduct ablation study and document the selection of hyperparameters in Section 4.4. Details about our experimental compute cost and complexity can be found in Appendix 5.
@@ 4.1 Experiment Setup
FLAN V2 (Longpre et al., 2023) is a widely-used instruction fine-tuning dataset that combines FLAN 2021, P3++, Super-Natural Instructions, along with additional reasoning, dialogue, and program synthesis datasets. We utilize a 99K subset of FLAN V2, processed by (Wang et al., 2023b), as our training data pool, from which we select prompts and annotations (i.e., responses).
@@
We conduct experiments with the 7B version of the prefix language model LLaMA-2 (Touvron et al., 2023), across different annotation budgets. Prior to fine-tuning, we choose a subset of prompts for annotation by either random sampling or using experimental design strategies, including uncertainty-based selection, k-Center selection, and submodular selections. These strategies are computed based on the prefix model only. Subsequently, we fine-tune the model on the annotated prompt/response pairs using Low-rank Adaptation (LoRA) (Hu et al., 2021).
@@
We adopt the similar evaluation tasks as in the original FLAN V2, using the MMLU (Hendrycks et al., 2020) and BBH (Suzgun et al., 2022) benchmarks to evaluate the zero-shot generalization capability of our fine-tuned model. Massive Multitask Language Understanding dataset (MMLU) is a classification task with a set of questions about 57 subjects ranging in difficulty from elementary levels to professional levels, broadly testing mode’s factual knowledge and reasoning. Big-Bench-Hard (BBH) is a generation task with 23 challenging tasks from Big-Bench (Srivastava et al., 2022), broadly testing models’ general reasoning capabilities. In our study, we specifically select 5-shot MMLU and a random 20% subset of Chain-Of-Thought BBH inputs, due to the computational resource limitations. In addition, we also evaluate our methods based on AlpacaEval (Li et al., 2023b), where given a prompt, the responses from two models are compared by GPT-4 turbo. The performance of the two models are reported as the win rate against each other.